# RAD: Training an End-to-End Driving Policy via Large-Scale 3DGS-based Reinforcement Learning

### v1

@@ Abstract
Existing end-to-end autonomous driving (AD) algorithms typically follow the Imitation Learning (IL) paradigm, which faces challenges such as causal confusion and the open-loop gap. In this work, we establish a 3DGS-based closed-loop Reinforcement Learning (RL) training paradigm. By leveraging 3DGS techniques, we construct a photorealistic digital replica of the real physical world, enabling the AD policy to extensively explore the state space and learn to handle out-of-distribution scenarios through large-scale trial and error. To enhance safety, we design specialized rewards that guide the policy to effectively respond to safety-critical events and understand real-world causal relationships. For better alignment with human driving behavior, IL is incorporated into RL training as a regularization term. We introduce a closed-loop evaluation benchmark consisting of diverse, previously unseen 3DGS environments. Compared to IL-based methods, RAD achieves stronger performance in most closed-loop metrics, especially 3×3$\times$3 × lower collision rate. Abundant closed-loop results are presented at https://hgao-cv.github.io/RAD.
@@ 1 Introduction
End-to-end autonomous driving (AD) is currently a trending topic in both academia and industry. It replaces a modularized pipeline with a holistic one by directly mapping sensory inputs to driving actions, offering advantages in system simplicity and generalization ability. Most existing end-to-end AD algorithms [11, 17, 49, 43, 22, 2, 38, 28] follow the Imitation Learning (IL) paradigm, which trains a neural network to mimic human driving behavior. However, despite their simplicity, IL-based methods face significant challenges in real-world deployment.
One key issue is causal confusion. IL trains a network to replicate human driving policies by learning from demonstrations. However, this paradigm primarily captures correlations rather than causal relationships between observations (states) and actions. As a result, IL-trained policies may struggle to identify the true causal factors behind planning decisions, leading to shortcut learning [8], e.g., merely extrapolating future trajectories from historical ones [23, 47]. Furthermore, since IL training data predominantly consists of common driving behaviors and does not adequately cover long-tailed distributions, IL-trained policies tend to converge to trivial solutions, lacking sufficient sensitivity to safety-critical events such as collisions.
@@
A straightforward solution to these problems is to perform closed-loop Reinforcement Learning (RL) training, which requires a driving environment that can interact with the AD policy. However, using real-world driving environments for closed-loop training poses prohibitive safety risks and operational costs. Simulated driving environments with sensor data simulation capabilities [6, 14] (which are required for end-to-end AD) are typically built on game engines [7, 39] but fail to provide realistic sensor simulation results.
In this work, we establish a 3DGS-based [18] closed-loop RL training paradigm. Leveraging 3DGS techniques, we construct a photorealistic digital replica of the real world, where the AD policy can extensively explore the state space and learn to handle out-of-distribution situations through large-scale trial and error. To ensure effective responses to safety-critical events and a better understanding of real-world causations, we design specialized safety-related rewards. However, RL training presents several critical challenges, which this paper addresses.
@@
To validate the effectiveness of our approach, we construct a closed-loop evaluation benchmark comprising diverse, unseen 3DGS environments. Our method, RAD, outperforms IL-based approaches across most closed-loop metrics, notably achieving a collision rate that is 3×3$\times$3 × lower.
@@
  • We combine RL and IL to synergistically optimize the AD policy. RL complements IL by modeling the causations and narrowing the open-loop gap, while IL complements RL in terms of human alignment.
  • We validate the effectiveness of RAD on a closed-loop evaluation benchmark consisting of diverse, unseen 3DGS environments. RAD achieves stronger performance in closed-loop evaluation, particularly a 3×3$\times$3 × lower collision rate, compared to IL-based methods.
@@ 2 Related Work
Dynamic Scene Reconstruction. Implicit neural representations have dominated novel view synthesis and dynamic scene reconstruction, with methods like UniSim [46], MARS [44], and NeuRAD [40] leveraging neural scene graphs to enable structured scene decomposition. However, these approaches rely on implicit representations, leading to slow rendering speeds that limit their practicality in real-time applications. In contrast, 3D Gaussian Splatting (3DGS) [18] has emerged as an efficient alternative, offering significantly faster rendering while maintaining high visual fidelity. Recent works have explored its potential for dynamic scene reconstruction, particularly in autonomous driving scenarios. StreetGaussians [45], DrivingGaussians [51], and HUGSIM [50] have demonstrated the effectiveness of Gaussian-based representations in modeling urban environments. These methods achieve superior rendering performance while maintaining controllability by explicitly decomposing scenes into structured components. However, these works primarily leverage 3DGS for closed-loop evaluation. In this work, we incorporate 3DGS into an RL training framework.
End-to-End Autonomous Driving. Learning-based planning has shown great potential recently due to its data-driven nature and impressive performance with increasing amounts of data. UniAD [12] demonstrates the potential of end-to-end autonomous driving by integrating multiple perception tasks to enhance planning performance. VAD [16] further explores the use of compact vectorized scene representations to improve efficiency. A series of works [20, 43, 49, 24, 42, 9, 3, 33] also adopt the single-trajectory planning paradigm and further enhance planning performance. VADv2 [2] shifts the paradigm towards multi-mode planning by modeling the probabilistic distribution of a planning vocabulary. Hydra-MDP [22] improves the scoring mechanism of VADv2 by introducing additional supervision from a rule-based scorer. SparseDrive [38] explores an alternative BEV-free solution. DiffusionDrive [28] proposes a truncated diffusion policy that denoises an anchored Gaussian distribution to a multi-mode driving action distribution. Most of the end-to-end methods follow the data-driven IL-training paradigm. In this work, we present an RL-training paradigm based on 3DGS.
Reinforcement Learning. Reinforcement Learning is a promising technique that has not been fully explored. AlphaGo [36] and AlphaGo Zero [37] have demonstrated the power of Reinforcement Learning in the game of Go. Recently, OpenAI O1 [32] and Deepseek-R1 [4] have leveraged Reinforcement Learning to develop reasoning abilities. Several studies have also applied Reinforcement Learning in autonomous driving [41, 1, 48, 31, 13]. However, these studies are based on non-photorealistic simulators (such as CARLA [6]) or do not involve end-to-end driving algorithms, as they require perfect perception results as input. To the best of our knowledge, RAD is the first work to train an end-to-end AD agent using Reinforcement Learning in a photorealistic 3DGS environment.
@@ 3.1 End-to-End Driving Policy
BEV Encoder. We first employ a BEV encoder [21] to transform multi-view image features from the perspective view to the Bird’s Eye View (BEV), obtaining a feature map in the BEV space. This feature map is then used to learn instance-level map features and agent features.
Map Head. Then we utilize a group of map tokens [27, 25, 26] to learn the vectorized map elements of the driving scene from the BEV feature map, including lane centerlines, lane dividers, road boundaries, arrows, traffic signals, etc.
Agent Head. Besides, a group of agent tokens [15] is adopted to predict the motion information of other traffic participants, including location, orientation, size, speed, and multi-mode future trajectories.
Image Encoder. Apart from the above instance-level map and agent tokens, we also use an individual image encoder [5, 10] to transform the original images into image tokens. These image tokens provide dense and rich scene information for planning, complementary to the instance-level tokens.
Figure 3: Post-training. $N$ workers parallelly run. The generated rollout data $(s_{t},a_{t},r_{t+1},s_{t+1},...)$ are recorded in a rollout buffer. Rollout data and human driving demonstrations are used in RL- and IL-training steps to fine-tune the AD policy synergistically.
Action Space. To accelerate the convergence of RL training, we design a decoupled discrete action representation. We divide the action into two independent components: lateral action and longitudinal action. The action space is constructed over a short $0.5$-second time horizon, during which the vehicle’s motion is approximated by assuming constant linear and angular velocities. Under this assumption, the lateral action $a^{x}$ and longitudinal action $a^{y}$ can be directly computed based on the current linear and angular velocities. By combining decoupling with a limited temporal scope and simplified motion model, our approach effectively reduces the dimensionality of the action space, accelerating training convergence.
@@
$\pi(a^{x}\mid s)=$ $\text{softmax}(\text{MLP}(\phi(E_{\text{plan}},E_{\text{scene}})$ (1)
$+E_{\text{navi}}+E_{\text{state}})),$
$\pi(a^{y}\mid s)=$ $\text{softmax}(\text{MLP}(\phi(E_{\text{plan}},E_{\text{scene}})$
$+E_{\text{navi}}+E_{\text{state}})),$
@@ 3.2 Training Paradigm
We select a large amount of risky dense-traffic clips from collected driving demonstrations. For each clip, we train an independent 3DGS model that reconstructs the clip and serves as a digital driving environment. As shown in Fig. 3, we set $N$ parallel workers. Each worker randomly samples a 3DGS environment and begins rollout, i.e., the AD policy controls the ego vehicle to move and iteratively interacts with the 3DGS environment. After the rollout process of this 3DGS environment ends, the generated rollout data $(s_{t},a_{t},r_{t+1},s_{t+1},...)$ are recorded in a rollout buffer, and the worker will sample a new 3DGS environment for another round of rollout.
As for policy optimization, we iteratively perform RL-training steps and IL-training steps. For RL-training steps, we sample data from the rollout buffer and follow the Proximal Policy Optimization (PPO) framework [35] to update the AD policy. For IL-training steps, we use real-world driving demonstrations to update the policy. After a fixed number of training steps, the updated AD policy is sent to every worker to replace the old one, to avoid a distribution shift between data collection and optimization. We only update the parameters of the image encoder and the planning head. The parameters of the BEV encoder, the map head, and the agent head are frozen. The detailed RL design is presented below.
@@ 3.3 Interaction Mechanism between AD Policy and 3DGS Environment
In the 3DGS environment, the ego vehicle acts according to the AD policy. Other traffic participants act according to real-world data in a log-replay manner. A simplified kinematic bicycle model is employed to iteratively update the ego vehicle’s pose at every $\Delta t$ seconds as follows:
$x_{t+1}^{w}$ $=x_{t}^{w}+v_{t}\cos\left(\psi_{t}^{w}\right)\Delta t,$ (3)
$y_{t+1}^{w}$ $=y_{t}^{w}+v_{t}\sin\left(\psi_{t}^{w}\right)\Delta t,$
$\psi_{t+1}^{w}$ $=\psi_{t}^{w}+\frac{v_{t}}{L}\tan\left(\delta_{t}\right)\Delta t,$
@@
Figure 4: Example diagram of four types of reward sources. (1): Collision with a dynamic obstacle ahead triggers a reward $r_{\text{dc}}$. (2): Hitting a static roadside obstacle incurs a reward $r_{\text{sc}}$. (3): Moving onto the curb exceeds the positional deviation threshold $d_{\text{max}}$, triggering a reward $r_{\text{pd}}$. (4): Drifting toward the adjacent lane exceeds the heading deviation threshold $\psi_{\text{max}}$, triggering a reward $r_{\text{hd}}$.
@@ 3.4 Reward Modeling
=$\mathcal{R}=\{r_{\text{dc}},r_{\text{sc}},r_{\text{pd}},r_{\text{hd}}\}.$ (4)
@@ 3.5 Policy Optimization
In the closed-loop environment, the error in each single step accumulates over time. The aforementioned rewards are not only caused by the current action but also by the actions of the preceding steps. The rewards are propagated forward with Generalized Advantage Estimation (GAE) [34] to optimize the action distribution of the preceding steps.
@@
$r_{t}^{x}$ $=r_{t}^{\text{sc}}+r_{t}^{\text{pd}}+r_{t}^{\text{hd}},$ (5)
$r_{t}^{y}$ $=r_{t}^{\text{dc}}.$
Similarly, the value function $V(s_{t})$ is decoupled into two components: $V_{x}(s_{t})$ for the lateral dimension and $V_{y}(s_{t})$ for the longitudinal dimension. These value functions estimate the expected cumulative rewards for the lateral and longitudinal actions, respectively. The advantage estimates $\hat{A}_{t}^{x}$ and $\hat{A}_{t}^{y}$ are then computed as follows:
$\delta_{t}^{x}$ $=r_{t}^{x}+\gamma V_{x}(s_{t+1})-V_{x}(s_{t}),$ (6)
$\delta_{t}^{y}$ $=r_{t}^{y}+\gamma V_{y}(s_{t+1})-V_{y}(s_{t}),$
$\hat{A}_{t}^{x}$ $=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{x},$
$\hat{A}_{t}^{y}$ $=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{y},$
To further clarify the relationship between the advantage estimates and the reward components, we decompose $\hat{A}_{t}^{x}$ and $\hat{A}_{t}^{y}$ based on the reward decomposition in Eq. 5 and the advantage estimation in Eq. 6. Specifically, we derive the following decomposition:
$\hat{A}_{t}^{x}$ $=\hat{A}_{t}^{\text{sc}}+\hat{A}_{t}^{\text{pd}}+\hat{A}_{t}^{\text{hd}},$ (7)
$\hat{A}_{t}^{y}$ $=\hat{A}_{t}^{\text{dc}},$
where $\hat{A}_{t}^{\text{sc}}$ is the advantage estimate for avoiding static collisions, $\hat{A}_{t}^{\text{pd}}$ is the advantage estimate for minimizing positional deviations, $\hat{A}_{t}^{\text{hd}}$ is the advantage estimate for minimizing heading deviations, and $\hat{A}_{t}^{\text{dc}}$ is the advantage estimate for avoiding dynamic collisions.
These advantage estimates are used to guide the update of the AD policy $\pi_{\theta}$, following the PPO framework [35]. By leveraging the decomposed advantage estimates $\hat{A}_{t}^{x}$ and $\hat{A}_{t}^{y}$, we can independently optimize the lateral and longitudinal dimensions of the policy. This is achieved by defining separate objective functions $\mathcal{L}_{x}^{\text{CLIP}}(\theta)$ and $\mathcal{L}_{y}^{\text{CLIP}}(\theta)$ for each dimension, as follows:
$\mathcal{L}^{\text{PPO}}(\theta)$ $=\mathcal{L}_{x}^{\text{PPO}}(\theta)+\mathcal{L}_{y}^{\text{PPO}}(\theta),$
The clipped objective function $\mathcal{L}^{\text{PPO}}(\theta)$ prevents excessively large updates to the policy parameters $\theta$, thereby maintaining training stability.

### v2

@@ Abstract
Existing end-to-end autonomous driving (AD) algorithms typically follow the Imitation Learning (IL) paradigm, which faces challenges such as causal confusion and an open-loop gap. In this work, we propose RAD, a 3DGS-based closed-loop Reinforcement Learning (RL) framework for end-to-end Autonomous Driving. By leveraging 3DGS techniques, we construct a photorealistic digital replica of the real physical world, enabling the AD policy to extensively explore the state space and learn to handle out-of-distribution scenarios through large-scale trial and error. To enhance safety, we design specialized rewards to guide the policy in effectively responding to safety-critical events and understanding real-world causal relationships. To better align with human driving behavior, we incorporate IL into RL training as a regularization term. We introduce a closed-loop evaluation benchmark consisting of diverse, previously unseen 3DGS environments. Compared to IL-based methods, RAD achieves stronger performance in most closed-loop metrics, particularly exhibiting a $3\times$ lower collision rate. Abundant closed-loop results are presented in the supplementary material. Code is available at https://github.com/hustvl/RAD for facilitating future research.
@@ 1 Introduction
End-to-end autonomous driving (AD) is currently a trending topic in both academia and industry. It replaces a modularized pipeline with a holistic one by directly mapping sensory inputs to driving actions, offering advantages in system simplicity and generalization ability. Most existing end-to-end AD algorithms uniad ; vad ; zheng2024genad ; paradrive ; hydramdp ; vadv2 ; sun2024sparsedrive ; diffusiondrive follow the Imitation Learning (IL) paradigm, which trains a neural network to mimic human driving behavior. However, despite their simplicity, IL-based methods face significant challenges in real-world deployment.
@@
One key issue is causal confusion. IL trains a network to replicate human driving policies by learning from demonstrations. However, this paradigm primarily captures correlations rather than causal relationships between observations (states) and actions. As a result, IL-trained policies may struggle to identify the true causal factors behind planning decisions, leading to shortcut learning shortcut , e.g., merely extrapolating future trajectories from historical ones egomlp ; admlp . Furthermore, since IL training data predominantly consists of common driving behaviors and does not adequately cover long-tailed distributions, IL-trained policies tend to converge to trivial solutions, lacking sufficient sensitivity to safety-critical events such as collisions.
@@
A straightforward solution to these problems is to perform closed-loop Reinforcement Learning (RL) training, which requires a driving environment that can interact with the AD policy. However, using real-world driving environments for closed-loop training poses prohibitive safety risks and operational costs. Simulated driving environments with sensor data simulation capabilities carla ; carsim (which are required for end-to-end AD) are typically built on game engines unreal ; unity but fail to provide realistic sensor simulation results.
In this work, we establish a 3DGS-based 3dgs closed-loop RL training paradigm. Leveraging 3DGS techniques, we construct a photorealistic digital replica of the real world, where the AD policy can extensively explore the state space and learn to handle out-of-distribution situations through large-scale trial and error. To ensure effective responses to safety-critical events and a better understanding of real-world causations, we design specialized safety-related rewards. However, RL training presents several critical challenges, which this paper addresses.
@@
To validate the effectiveness of our approach, we construct a closed-loop evaluation benchmark comprising diverse, unseen 3DGS environments. Our method, RAD, outperforms IL-based approaches across most closed-loop metrics, notably achieving a collision rate that is $3\times$ lower.
@@
  • We propose to combine RL and IL to synergistically optimize the end-to-end AD policy. RL complements IL by modeling the causations and narrowing the open-loop gap, while IL complements RL in terms of human alignment.
  • We validate the effectiveness of RAD on a closed-loop evaluation benchmark consisting of diverse, unseen 3DGS environments. RAD achieves stronger performance in closed-loop evaluation, particularly a $3\times$ lower collision rate, compared to IL-based methods.
@@ 2 Related Work
Dynamic Scene Reconstruction. Implicit neural representations have been widely used in novel view synthesis and dynamic scene reconstruction, as in UniSim UniSim , MARS MARS , and NeuRAD NeuRAD , which leverage neural scene graphs for structured decomposition. However, their slow rendering speeds hinder real-time applications. Several recent works StreetGaussians ; DrivingGaussian ; hugsim have demonstrated the effectiveness of 3D Gaussian Splatting (3DGS) 3dgs for dynamic urban scene reconstruction. While prior works UniSim ; MARS ; hugsim primarily utilize reconstructed scenes for closed-loop evaluation, we go a step further by incorporating scenes reconstructed via 3DGS into the RL training loop to enhance policy learning.
End-to-End Imitation Learning for Autonomous Driving. Recent advances in learning-based planning have shown strong potential, driven by large-scale data. UniAD uniad integrates multiple perception tasks to boost planning, while VAD vad improves efficiency with compact vectorized representations. Follow-up works li2024enhancing ; paradrive ; zheng2024genad ; egomlp ; wang2024driving ; gu2024producing ; chen2025ppad ; transfuser ; sun2024sparsedrive enhance the single-trajectory paradigm, whereas VADv2 vadv2 and Hydra-MDP hydramdp shift towards multi-modal planning with improved scoring. DiffusionDrive diffusiondrive proposes a truncated diffusion policy that denoises an anchored Gaussian distribution to a multi-mode driving action distribution. While existing methods largely follow an IL paradigm, we enhance it by incorporating closed-loop RL to further optimize the policy.
Reinforcement Learning. Reinforcement Learning is a promising technique that has not been fully explored. AlphaGo alphago and AlphaGo Zero alphago-zero have demonstrated the power of Reinforcement Learning in the game of Go. Recently, OpenAI O1 openai-o1 and Deepseek-R1 deepseek-r1 have leveraged Reinforcement Learning to develop reasoning abilities. RL has also been applied in autonomous driving toromanoff2020end ; chen2021learning ; roach ; ILNotEnough ; GUMP , though often using non-photorealistic simulators (e.g., CARLA carla ) or requiring perfect perception inputs. To the best of our knowledge, RAD is the first to train an end-to-end autonomous driving agent using RL in a photorealistic 3DGS environment.
Combining Imitation and Reinforcement Learning. To combine the sample efficiency of IL with the robustness of RL, CADRE zhao2022cadre and CIRL liang2018cirl adopt a two-stage pipeline, first imitating expert demonstrations and then fine-tuning policies with RL to improve success rates in basic and complex CARLA scenarios. Lu et al. ILNotEnough and Huang et al. huang2022efficient instead perform joint IL and RL optimization, where IL constrains RL exploration to maintain human-like behaviors while improving robustness and rare-case success. Although these approaches demonstrate the benefits of combining IL and RL, they are still limited to non-photorealistic simulators or rely on structured BEV representations and virtual sensor data. In contrast, RAD is the first to jointly integrate IL and RL within a 3DGS-based photorealistic digital-twin environment, enabling fully end-to-end policy learning from real-world sensor inputs.
@@ 3.1 End-to-End Driving Policy
BEV Encoder. We first employ a BEV encoder li2022bevformer to transform multi-view image features from the perspective view to the Bird’s Eye View (BEV), obtaining a feature map in the BEV space. This feature map is then used to learn instance-level map features and agent features.
Map Head. Then we utilize a group of map tokens maptrv2 ; liao2022maptr ; lanegap to learn the vectorized map elements of the driving scene from the BEV feature map, including lane centerlines, lane dividers, road boundaries, arrows, traffic signals, etc.
Agent Head. Besides, a group of agent tokens jiang2022pip is adopted to predict the motion information of other traffic participants, including location, orientation, size, speed, and multi-mode future trajectories.
Image Encoder. Apart from the above instance-level map and agent tokens, we use an individual image encoder vit ; he2016resnet to transform the original images into image tokens. These image tokens provide dense and rich scene information for planning, complementary to the instance-level tokens.
Action Space. We propose a decoupled discrete action representation by separating lateral and longitudinal actions. Each action is defined over a short 0.5-second horizon, assuming constant linear and angular velocities. This modeling assumption allows direct computation of control signals from the current action predictions. The resulting formulation reduces the dimensionality of the action space, facilitating more efficient policy optimization.
@@
$\pi(a^{x}\mid s)=$ $\text{softmax}(\text{MLP}(\phi(E_{\text{plan}},E_{\text{scene}})$ $+E_{\text{navi}}+E_{\text{state}})),$ (1)
$\pi(a^{y}\mid s)=$ $\text{softmax}(\text{MLP}(\phi(E_{\text{plan}},E_{\text{scene}})$ $+E_{\text{navi}}+E_{\text{state}})),$
@@ 3.2 Training Paradigm
[…]
For policy optimization, we alternate between RL and IL steps. In RL steps, the Proximal Policy Optimization (PPO) PPO is used to update the policy from the rollout buffer. In IL steps, expert demonstrations are used for supervised updates. After a fixed number of steps, the updated policy is synchronized across all workers to mitigate distributional shift. During training, only the image encoder and planning head are updated, while the BEV encoder, map head, and agent head are kept frozen. The detailed RL design is introduced below.
@@ 3.3 Interaction Mechanism between AD Policy and 3DGS Environment
In the 3DGS environment, the ego vehicle acts according to the AD policy. Other traffic participants can be controlled in many manners: 1) using a smart agent model; 2) using Intelligent Driver Model (IDM) for route tracking; 3) log-replay with real-world data. In order to fully recover the real-world traffic flow, we use log-replay to control other traffic participants in experiments. A conventional kinematic bicycle model is employed to iteratively update the ego vehicle’s pose at every $\Delta t$ seconds as follows:
$x_{t+1}^{w}=x_{t}^{w}+v_{t}\cos\left(\psi_{t}^{w}\right)\Delta t,\;y_{t+1}^{w}=y_{t}^{w}+v_{t}\sin\left(\psi_{t}^{w}\right)\Delta t,\;\psi_{t+1}^{w}=\psi_{t}^{w}+\frac{v_{t}}{L}\tan\left(\delta_{t}\right)\Delta t,$ (3)
@@
Table 1: Comparison of training strategies in Stage 3.
[…]
$\downarrow Jerk\downarrow$
IL 0.229 0.211 0.018 0.066 0.039 0.027 0.238 3.928 0.103
RL 0.143 0.128 0.015 0.080 0.065 0.015 0.345 4.204 0.085
RL+IL 0.089 0.080 0.009 0.063 0.042 0.021 0.257 4.495 0.082
Figure 4: Example diagram of four types of reward sources. (1) Collision with a dynamic obstacle ahead triggers $r_{\text{dc}}$. (2) Hitting a static roadside obstacle incurs $r_{\text{sc}}$. (3) Moving onto the curb triggers $r_{\text{pd}}$. (4) Drifting toward the adjacent lane triggers $r_{\text{hd}}$.
@@ 3.4 Reward Modeling
$\mathcal{R}=\{r_{\text{dc}},r_{\text{sc}},r_{\text{pd}},r_{\text{hd}}\}.$ (4)
@@ 3.5 Policy Optimization
In the closed-loop environment, the error in each single step accumulates over time. The aforementioned rewards are not only caused by the current action but also by the actions of the preceding steps. The rewards are propagated forward with Generalized Advantage Estimation (GAE) gae to optimize the action distribution of the preceding steps.
@@
$r_{t}^{x}=r_{t}^{\text{sc}}+r_{t}^{\text{pd}}+r_{t}^{\text{hd}},\quad r_{t}^{y}=r_{t}^{\text{dc}}.$ (5)
Similarly, the value function $V(s_{t})$ is decoupled into two components: $V_{x}(s_{t})$ for the lateral dimension and $V_{y}(s_{t})$ for the longitudinal dimension. These value functions estimate the expected cumulative rewards for the lateral and longitudinal actions, respectively. The advantage estimates $\hat{A}_{t}^{x}$ and $\hat{A}_{t}^{y}$ are computed using GAE (detailed in Appendix A.3), then decomposed into components aligned with reward types:
$\hat{A}_{t}^{x}=\hat{A}_{t}^{\text{sc}}+\hat{A}_{t}^{\text{pd}}+\hat{A}_{t}^{\text{hd}},\quad\hat{A}_{t}^{y}=\hat{A}_{t}^{\text{dc}}.$ (6)
The policy $\pi_{\theta}$ is optimized via a modified PPO objective:
$\mathcal{L}^{\text{PPO}}(\theta)=\mathcal{L}_{x}^{\text{PPO}}(\theta)+\mathcal{L}_{y}^{\text{PPO}}(\theta),$ (7)
with independent clipping mechanisms for each dimension (full equations in Appendix A.3). The clipped objective prevents excessively large updates to the policy parameters $\theta$.
3.6 Auxiliary Objective
To address the sparse reward challenge in RL, we design directional auxiliary objectives that provide dense supervision for both longitudinal and lateral controls. These objectives adaptively adjust action probabilities based on real-time collision risk and trajectory deviation, offering more informative gradients during training.
Specifically, a dynamic collision auxiliary loss $\mathcal{L}_{\text{dc}}$ penalizes acceleration and encourages deceleration when past collisions with preceding vehicles are detected. A static collision auxiliary loss $\mathcal{L}_{\text{sc}}$ suppresses steering toward static obstacles and promotes avoidance behavior. For trajectory alignment, the positional deviation loss $\mathcal{L}_{\text{pd}}$ encourages lateral corrections toward the reference path, while the heading deviation loss $\mathcal{L}_{\text{hd}}$ promotes angular corrections to minimize orientation error. Additional implementation details are provided in Appendix A.4.
The composite optimization objective integrates the PPO policy gradient with auxiliary guidance:
$\mathcal{L}(\theta)=\mathcal{L}^{\text{PPO}}(\theta)+\lambda_{1}\mathcal{L}_{\text{dc}}(\theta)+\lambda_{2}\mathcal{L}_{\text{sc}}(\theta)+\lambda_{3}\mathcal{L}_{\text{pd}}(\theta)+\lambda_{4}\mathcal{L}_{\text{hd}}(\theta),$ (8)
Table 2: Impact of different reward components on performance.